OpenStudy (anonymous):

Help!! Medal!! Determine which ordered pair in the form (x, y) is a solution to 3x – 4y = 21. a. (-2, -3) b. (0, 7) c. (7, 0) d. (-3, 2)

OpenStudy (anonymous):

All you have to do is Plug in the numbers. 3(-2) - 4(-3) does that equal 21?

OpenStudy (anonymous):

@pigglet

OpenStudy (anonymous):

no? i got distracted, sorry

OpenStudy (anonymous):

It's okay. So, you know the first one isn't right. Now do the second one.

OpenStudy (anonymous):

3*0-4*7?

OpenStudy (anonymous):

That is the correct step up now what is the answer to answer choice B?

OpenStudy (anonymous):

-28?

OpenStudy (anonymous):

@pigglet

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):

3(0) - 4(7) 3 * 0 = 0 4 * 7 = 28 28 + 0 = 28 and C is correct